# Optimal Data Splitting for Holdout Cross-Validation in Large Covariance Matrix Estimation

### v1

@@ 3.2.1 Optimal split
The leading term of the asymptotic expansion of equation (54), for $n\to\infty$ simplifies to
@@
Then the holdout error formula of equation (45) converges to
@@
which coincides with equation (22), and proves the convergence to the expected Frobenius error of the oracle estimator. ∎
@@
It is worth remarking that the minimum in the right panel of figure 2, around $k=1.5$, is a prediction confirmed by our formula written in equation (54). Such a minimum is counterintuitive since for $k=1.5$, the prediction of the formula implies a test set double the train set, which is against every recommendation for CV.
@@ 4 Conclusion
In this work, we computed analytically the estimation error as expected Frobenius error of the holdout estimator for Gaussian stationary data. We obtained a closed-form expression for the case of white inverse Wishart population matrices, while for the general case numerical approximations are needed. Our approach involved the use of Wick’s theorem to link the Frobenius error of estimation with the moments of the population covariance matrix and then the recognition of the oracle eigenvalues. Furthermore, the method developed by Ledoit and Péché for computing the oracle eigenvalues can be applied to a wider range of population covariance matrix distributions beyond the inverse Wishart case. This suggests that our approach could be extended to derive analytical formulas for the holdout error in these more general settings, opening up interesting avenues for future research.
With our approach, we found that the optimal split that minimizes the estimation error is proportional to the square root of the matrix dimension. The appearance here of a power law should be further investigated in future works as they often hold significance for applications and one can wonder whether a similar pattern would remain valid for other noise models and population covariance distributions.

### v2

@@ 3.2.1 Optimal split
The leading term of the asymptotic expansion of Eq. (56), for $n\to\infty$ simplifies to
@@
Then the holdout error formula of Eq. (47) converges to
@@
which coincides with Eq. (22), and proves the convergence to the expected Frobenius error of the oracle estimator. ∎
This result was expected considering Lam (2016, 2015) convergence result written in therorem 2.2, but in our setting we can be a bit less restrictive about the choice of the split in the high-dimensional regime. For example for a $k$ proportional to $n^{2/3}$, we would obtain a $t_{\textrm{out}}$ proportional to $n^{1/3}$ which would not verify the convergence assumption of Lam but still verifies the assumptions of the previous corollary.
@@
It is worth remarking that the minimum in the right panel of Fig. 2, around $k=1.5$, is a prediction confirmed by our formula written in Eq. (56). Such a minimum is counterintuitive since for $k=1.5$, the prediction of the formula implies a test set double the train set.
@@ 4 Conclusion
In this work, we computed analytically the expected Frobenius error of the holdout estimator for Gaussian stationary data. We obtained a closed-form expression for the case of white inverse Wishart population matrices, while for the general case numerical approximations are needed. Our approach involved the use of Wick’s theorem to link the Frobenius error of estimation with the moments of the population covariance matrix and then the recognition of the oracle eigenvalues. Furthermore, the method developed by Ledoit and Péché for computing the oracle eigenvalues can be applied to a wider range of population covariance matrix distributions beyond the inverse Wishart case. This suggests that our approach could be extended to derive analytical formulas for the holdout error in these more general settings, opening up interesting avenues for future research.
With our approach, we found that the optimal split that minimizes the estimation error is proportional to the square root of the matrix dimension. The appearance here of a power law should be further investigated in future works as they often hold significance for applications and one can wonder whether a similar pattern would remain valid for other noise models and population covariance distributions. In the finite but large limit, the error sharpens around one optimal split instead of the plateau obtained in the asymptotic regime, making the choice of the split particularly important. This paves the way for further theoretical studies that could in the future help practitioners calibrate the holdout or CV estimators in the finite-sample case instead of fitting the split empirically directly on the data.